haemoglobin contains 0.334% of iron by weight. the molecular weight of haemoglobin in approximately 67200. the number of iron atoms present in one molecule of haemoglobin is ? (atomic weight of iron is 56) 
 

Grade:12

4 Answers

sohini paul
7 years ago
100 g haemoglobin contains 0.334 g iron
then, 67200 g haemoglobin contains =224.44 g iron.
 
Now , 56 g iron contains=6.023 * pow(10,23) atoms of iron
then 244.44g contains =26.3*pow(10,23) atoms of iron.
 
now , 6.022*pow(10,23) molecules of haemoglobin=26.3*pow(10,23) atoms of iron
(since 67200g 0f haemoglobin =6.022*pow(10,23)moleculesof haemoglobin)
then, 1 molecule of haemoglobin contains= 4atoms of iron .
Dheeraj jareda
6 years ago
Percentage of Fe = 0.33% molecular weight of haemoglobinWeight of Fe present = (0.33/100) × 67200 = 221.8one Fe weighs 55.8 gTherefore, 221.8/55.8 = 4Hence, 4 atoms of iron are present in one molecule of haemoglobin.
Shubham
6 years ago
%by mass of Fe = mass of Fe atoms/molmass×1000.334=x×56/67200×100X=0.334×67200/56×100X=4No. Of Fe atoms are 4
ankit singh
3 years ago
then, 1 molecule of haemoglobin contains= 4atoms of iron ." then, 67200 g haemoglobin contains =224.44 g iron. then 244.44g contains =26.3*pow(10,23) atoms of iron.